OpenStudy (anonymous):

given that a^1/2 - a^-1/2=1,prove that a + a^-1=3

OpenStudy (eric_d):

a-2+a^-1=1

OpenStudy (eric_d):

thus, a+a^-1=3

OpenStudy (eric_d):

am i rite

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

\[\sqrt a-\frac{1}{\sqrt a}=1~~\iff~~a-1=\sqrt a~~\iff~~a^2-3a+1=0\] Finding the roots, you should have \[a=\frac{3\pm\sqrt{5}}{2}\] Show that \[a+\frac{1}{a}=\frac{3\pm\sqrt{5}}{2}+\frac{2}{3\pm\sqrt5}=3\]